OpenStudy (anonymous):

What is the exact circumference of a circle with a diameter of 17 inches? Possible answers: 17p inches 34p inches 8.5p inches 72.25p inches 289p inches I keep getting 53.4..what am I doing wrong?

OpenStudy (diyadiya):

\[Circumference = 2 \times \pi \times radius\]\[=2 \times \pi \times \frac{17}{2}\]\[=17 \pi\ inches\]

OpenStudy (anonymous):

I got that using a calculator but disregarded it for whatever reason..meh. Thanks a lot.

OpenStudy (diyadiya):

Haha You're Welcome :)